OpenStudy (anonymous):

(4x)/2y multiplied (3y)/8xy

OpenStudy (anonymous):

\(\dfrac{4x}{2y} \times \dfrac{3y}{8xy}\) Multiply straight across. Numerator times numerator. Denominator times denominator.

OpenStudy (cggurumanjunath):